OpenStudy (anonymous):

What is the solution of the following system? x - y=11 -x + y= -11

OpenStudy (anonymous):

there is an infinite amount of solutions

OpenStudy (anonymous):

Do you want a proof or do you know why

OpenStudy (anonymous):

x = y + 11 -x + y= -11 -(y+11) + y = -11 -y - 11 + y = -11 y - 11 = y - 11 y = y when you have a number equal to itself ex 0=0, then there is an infinite amount of solutions. y could be anything for it to be a solution